OpenStudy (anonymous):

What is the sum of the binary numbers? 111 + 100 (Points : 1) 1011 1111 1110 1101

OpenStudy (anonymous):

\[111_2=2^2+2^1+2^0=4+2+1=7_{10}\] \[100_2=2^2=4_{10}\] So \(111_2+100_2=11_{10}\). Convert 11 to base 2.